OpenStudy (thadds2003):

Solve the equation. x - 2/7 = 3 A. x=-3 2/7 B. x=-2 5/7 C. x= 2 5/7 D. x= 3 2/7

OpenStudy (calculusxy):

You want to isolate x and solve for it. To do so, you need to get rid of the constants first and in this case it would be -2/7. Since this is a subtraction, \(x \large \color{red}{-}\frac{2}{7}\) you would need to add \(\frac{2}{7}\) to both sides of the equal sign to make sure that there is a balance. Not only that but you will also cancel -2/7 out and have x alone (-2/7 + 2/7 = 0).

OpenStudy (calculusxy):

\[x - \frac{2}{7} = 3 \rightarrow \color{green}{x - \frac{2}{7} + \frac{2}{7} = 3 + \frac{2}{7} \rightarrow x = ? }\]

OpenStudy (calculusxy):

What will x equal to?

OpenStudy (thadds2003):

I dont get it......

OpenStudy (thadds2003):

@Anaise

OpenStudy (calculusxy):

you need to add 3 + 2/7